OpenStudy (anonymous):

The basketball team scored 12 points the first game and 47 points the eighth game of the season. How many points will they earn with this pattern on the 14th game of the season?

OpenStudy (anonymous):

(1, 12) (8, 47) Find slope: 47 - 12 = 35 8 - 1 = 7 slope = 35/7 = 5/1 6 games left: 5 points each 5*6 = 30 30 + 47 = 77 points Bump it so someone else can check; I'm not entirely sure about this answer.

OpenStudy (anonymous):

thank you i got 77 also but wasn't for sure